OpenStudy (anonymous):

what is 10 divided by 2/5

OpenStudy (twizttiez):

1

OpenStudy (anonymous):

ok thank you

OpenStudy (kamibug):

Use KCF (Keep, Change, Flip) 10/1 x 5/2 = 50/2 = 25

OpenStudy (twizttiez):

Wait its 25 oops

OpenStudy (twizttiez):

Sorry lol

OpenStudy (anonymous):

thx!!!

OpenStudy (twizttiez):

I do apologize pheonix XD

OpenStudy (anonymous):

lol

OpenStudy (isaiah.feynman):